OpenStudy (anonymous):

Use the definition of logarithm to simplify each expression log10 10

OpenStudy (anonymous):

is that \( log_{10}(10) \) or \(log(10) \cdot 10\) or what?

OpenStudy (anonymous):

log\[\log_{10} 10\]

OpenStudy (anonymous):

ITS EQUAL TO 1

OpenStudy (anonymous):

The definition of \(log_bx\) is that \[ b^y = x \] Therefore, \( log_{10}(10) \) is actually \[ 10^y = 10 \] Solve for 'y', that is the value of the expression \( log_{10}(10) \)